# XTransfer: Cross-Modality Model Transfer for Human Sensing with Few Data at the Edge

### v1

@@ 6.2 XTransfer performance
Repair performance. We first evaluate how our Repair loss reduces MMC shift to enhance the S-score of each layer with different MMC dimensionalities in the SRR pipeline. We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting. The channel removal is disabled in this experiment. We test with 4 different levels of MMC dimensionality (i.e., number of channels) ranging from 64 to 512. We employ N-Pair loss (Sohn2016, ) and Triplet loss (Schroff2015, ) as baselines. We run 15 rounds for each setting by default throughout the evaluation. Fig. 4(a) presents that our Repair loss consistently achieves the highest S-score of 0.2 on average, outperforming both N-Pair loss (0.05) and Triplet loss (0.07), across all levels of MMC dimensionality. It indicates that Repair loss effectively enhances layer performance.
@@
Efficient search performance. We now evaluate LWS control performance regarding model accuracy and layer search efficiency. We employ HHAR as target, ResNet18 on miniImageNet and Multi-ResNet18 on 5 image datasets as sources shown in Tab. 2, with a default search depth of 3 in 3-, 5- and 10-shot settings. We also utilize MetaSense in Tab. 1 as the oracle baseline. We compare different configurations: ResNet18 (Single), Multi-ResNet18 (Multi) without efficient search enabled, and Multi-ResNet18 (Multi-Efficient) with efficient search enabled. As expected, Fig. 4(b) shows that Multi-Efficient achieves a substantial time reduction ranging from 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline. Additionally, Multi-Efficient improves average accuracy by 7.8% over Multi across the three settings. While Multi-Efficient takes a longer search time than Single, it successfully achieves superior accuracy, highlighting the benefits of using multiple source models in XTransfer. The results also reveal that both Multi-Efficient and Multi achieve lower standard deviation on average compared to Single, suggesting that using multiple source models offers commendable search stability. In short, the results prove that the proposed efficient search significantly accelerates the layer search process and achieves better model accuracy.
@@
Impact of different sources. Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models. We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Tab. 2. We keep the other experimental setups unchanged. Fig. 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average, compared to that of 69.6%, 64.3%, and 67% achieved by the models using Text, Audio, Sensing as sources, respectively. This suggests that the latent features learned from Image have better discriminability than those from other modalities. Since XTransfer prioritizes accuracy, we select Image source models by default. More ablation study results refer to Appendix C.
@@ 6.3 Comprehensive comparison
Cross-modality model transfer performance. We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets, as shown in Tab. 2. We compare XTransfer against a range of state-of-the-art baselines shown in Tab. 1. We also mark our methods as Our-Single and Our-Multi, using pre-trained ResNet18 on miniImageNet and Multi-ResNet18 with 5 corresponding image datasets as sources. In addition, we evaluate the fine-tuning performance of GPT2-small for comparison and align the parameters with the work (zhou2023one, ). All methods on target datasets are evaluated under the same cross-modality FSL settings (see App. A), with MetaSense (metasense, ) serving as the oracle baseline, trained on Conv5 with sufficient target sensor data (Tab. 3). For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels. As shown in Tab.4 and 5, both our methods outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings. Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively. Interestingly, in 3-shot setting, the averaged model accuracy of our methods is slightly lower than that of the oracle on the HHAR, Gesture, and Emotion datasets. This suggests that XTransfer’s Repair loss may lead to less stable improvements for latent feature alignment when using extremely low sample sizes, but its performance remains strong and stable in 5- and 10-shot settings. In summary, the comprehensive results show that XTransfer achieves state-of-the-art model accuracy, substantially reducing the cost of data collection and model resource overhead for scalable deployment of human sensing applications.
@@
Training resource usage. We also examine the training resource usage of XTransfer by measuring GPU utilization (Util) on the cloud. We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Finn2017, ) and DAPN (Zhao2021, )), excluding methods that rely solely on fine-tuning or distance-based techniques. The experimental setup remains the same as above. As shown in Tab. 6, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines. Notably, both Our-Single and Our-Multi achieve a substantial reduction in GPU Util by 2.6 to 3.6 times and 2.5 to 3.5 times on average, respectively, compared to baselines. This is due to our layer-wise method, which progressively repairs and recombines only a subset of layers rather than training the entire model at once. It also highlights the effectiveness of our compact connector design. Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency. Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer with low computational overhead.
@@
MetaSense (metasense, ) 63.2 69.0 75.2 55.1 64.0 68.3 66.3 73.4 79.4 72.1 83.3 89.6 55.5 56.3 65.5 24.8 28.1 31.7
ProtoNet (Snell2017, ) 50.6 45.7 49.8 43.9 49.3 52.0 50.8 55.8 59.2 69.5 78.7 73.4 51.3 49.2 54.8 21.7 23.4 24.5
DAPN (Zhao2021, ) 48.7 51.2 54.7 56.9 61.2 63.1 45.8 49.4 50.5 75.1 78.6 79.6 48.0 50.0 58.7 24.0 24.4 25.5
MAML (Finn2017, ) 36.6 42.3 42.4 39.6 57.9 65.6 37.2 41.3 39.4 31.6 39.7 35.2 22.3 26.9 31.1 22.5 20.4 21.9
GPT2 (zhou2023one, ) 41.0 34.0 45.0 48.0 35.0 36.0 49.0 54.0 63.0 52.0 71.0 75.0 - - - - - -
@@
ProtoNet (Snell2017, ) 0.51 0.46 0.50 0.44 0.49 0.52 0.51 0.56 0.59 0.69 0.79 0.73 0.51 0.49 0.55 0.22 0.23 0.25
DAPN (Zhao2021, ) 0.49 0.51 0.55 0.57 0.61 0.63 0.46 0.49 0.51 0.75 0.79 0.80 0.48 0.50 0.59 0.24 0.24 0.26
MAML (Finn2017, ) 0.37 0.42 0.42 0.40 0.58 0.66 0.37 0.41 0.39 0.32 0.40 0.35 0.22 0.27 0.31 0.22 0.20 0.22
GPT2 (zhou2023one, ) 0.11 0.09 0.12 0.13 0.09 0.09 0.13 0.14 0.17 0.14 0.19 0.20 - - - - - -
@@
DAPN (Zhao2021, ) 48.85 49.19 51.79 44.62 49.49 48.81 42.73 48.71 50.59 43.88 56.63 57.51 44.67 59.73 63.12 46.66 47.32 48.48
MAML (Finn2017, ) 59.59 79.32 89.72 52.58 70.27 77.52 53.11 66.75 73.21 72.56 85.47 92.22 63.39 72.78 75.01 44.52 55.46 61.36
@@ 7 Conclusion
This paper presents XTransfer, a novel method that enables resource-efficient, modality-agnostic model transfer for human sensing applications at the edge. XTransfer harnesses the power of pre-trained models as free sources and requires only few sensor data to create compact models. It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a practical and scalable method for advancing human sensing.
Limitations and future work. Since XTransfer reuses layers from one or more pre-trained source models, its performance depends on their quality (c.f. Sec. 4). While performance varies with different source models, XTransfer’s modality-agnostic design ensures broad adaptability, enabling flexible use of diverse source models without concern for modality shift or resource constraints. For fair benchmarking, we primarily use the pre-trained ResNet18 from PyTorch Hub (PyTorchhub, ) (c.f. Sec. 6.1), but XTransfer’s layer-wise repairing and recombining support flexible backbone choices. While current version requires source models with the same backbone, future versions will support mixed backbones to further enhance flexibility.
@@ A.1 Settings for learning with few data
A typical FSL method consists of two stages–meta-training and meta-testing [56]. In meta-training stage, FSL requires a large-scale source dataset from the same singular modality (i.e., typically specific to the application) to train source models from scratch. The source dataset is divided into a support set and a query set  [^4] The support set and query set are assumed to be drawn from the same data distribution [82].. Data for each set is loaded according to the n-way (i.e., number of classes) and n-shot (i.e., number of samples per class) format [15]. The meta-testing stage follows the same structure but uses unseen target data. Specifically, the target support set is used to adapt the source models, which are then evaluated on the target query set. Leave-One-Out Cross-Validation (LOOCV) is widely used as the de facto standard for evaluating datasets involving multiple users in human sensing [18]. LOOCV works by leaving one user out as the target for meta-testing while using the remaining users to train the source models during meta-training in each round. The MetaSense [18] is used as an oracle baseline under FSL settings, where sufficient target sensing datasets are provided during meta-training stage. Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same singular modality [23]. In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but with pre-trained source models selected across different modalities without constraints.
@@ A.2 Setups in preliminary study
As shown in Fig. 2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ [51]. In Fig. 2(b), we test 3 source models (using a pre-trained ResNet18 from PyTorch Hub [38]) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Tab. 2. We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Tab. 2. We use a default 5-shot cross-modality FSL setting with LOOCV enabled.
@@ A.3 Setups in layer-wise analysis
In Section 3, we conduct a layer-wise analysis to gain insights into cross-modality model transfer with few sensor data. To set up, we continue to use the standard ResNet18 [38] and segment it as 9 L-Blocks, as shown in Tab. 3. We also employ the default reshaping (c.f. App. B.1) on HHAR as target with 5-shot setting, and use the original source model input (miniImageNet) as reference, marked as Original. We examine the single-source baselines including Transfer Learning (TL) [68], Fine-tuning (FT) [23], MAML [70], DAPN [52], and MetaSense [18] as the oracle baseline, shown in Tab. 1. To precisely track the MMC shift changes at each layer during fine-tuning,
@@ A.4 Layer-wise metric
To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) [63] as the layer-wise metric to evaluate how well the features are being fine-tuned. The S-score is in the range from -1 to 1 and being better if closing to 1. To verify whether the S-score indicates the layer performance, we observe the changes in the S-score and accuracy at each layer using the TL [68] and Original methods. Fig. 6(a) also shows that S-score and accuracy by Original are positively correlated, converging in a layer-wise manner. TL’s S-score also performs equivalently to its accuracy (i.e., S-score drop along with accuracy drop) in detecting the negative transfer across different layers.
@@ A.5 Implementation
We implement XTransfer using PyTorch version 1.4.0. We also benchmark XTransfer with a full development cycle of cross-modality FSL tasks, including pre- and post-processing, training, and inference on commercial edge devices. Once pre-trained source models and few sensor data are processed in cloud via XTransfer, compact models can be seamlessly converted to various DL platforms (e.g., PyTorch Mobile, TFLite) for broader edge deployments. To enable reproducibility of the results, we fix all random seeds in the code, set the evaluation mode consistently when evaluating output models, and enable CUDA convolution determinism. To measure model resource overhead (e.g., FLOPs and parameters), we use the standard PyTorch OpCounter. For regression tasks, class-based transformation [83] enables compatibility.

### v2

@@ 6.2 XTransfer performance
Repair performance. We first evaluate how our Repair loss reduces MMC shift to enhance the S-score of each layer with different MMC dimensionalities in the SRR pipeline. We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting. The channel removal is disabled in this experiment. We test with 4 different levels of MMC dimensionality (i.e., number of channels) ranging from 64 to 512. We employ N-Pair loss (Zhang et al., 2023c) and Triplet loss (Weller et al., 2022) in FSL sensing tasks as baselines. We run 15 rounds for each setting by default throughout the evaluation. Figure 4(a) presents that our Repair loss consistently achieves the highest S-score of 0.2 on average, outperforming both N-Pair loss (0.05) and Triplet loss (0.07), across all levels of MMC dimensionality. It indicates that Repair loss effectively enhances layer performance.
@@
Efficient search performance. We now evaluate LWS control performance regarding model accuracy and layer search efficiency. We employ HHAR as target, ResNet18 on miniImageNet and Multi-ResNet18 on 5 image datasets as sources shown in Table 2, with a default search depth of 3 in 3-, 5- and 10-shot settings. We also utilize MetaSense in Table 1 as the oracle baseline. We compare different configurations: ResNet18 (Single), Multi-ResNet18 (Multi) without efficient search enabled, and Multi-ResNet18 (Multi-Efficient) with efficient search enabled. As expected, Figure 4(b) shows that Multi-Efficient achieves a substantial time reduction ranging from 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline. Additionally, Multi-Efficient improves average accuracy by 7.8% over Multi across the three settings. While Multi-Efficient takes a longer search time than Single, it successfully achieves superior accuracy, highlighting the benefits of using multiple source models in XTransfer. The results also reveal that both Multi-Efficient and Multi achieve lower standard deviation on average compared to Single, suggesting that using multiple source models offers commendable search stability. In short, the results prove that the proposed efficient search significantly accelerates the layer search process and achieves better model accuracy.
@@
Impact of different sources. Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models. We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Table 2. We keep the other experimental setups unchanged. Figure 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average, compared to that of 69.6%, 64.3%, and 67% achieved by the models using Text, Audio, Sensing as sources, respectively. This suggests that the latent features learned from Image have better discriminability than those from other modalities. Since XTransfer prioritizes accuracy, we select Image source models by default. More ablation study results refer to Appendix C.
@@ 6.3 Comprehensive comparison
Cross-modality FSL performance. We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets (Table 2). We compare XTransfer against a range of SOTA baselines shown in Table 1. We also mark our methods as Our-Single and Our-Multi, using pre-trained ResNet18 on miniImageNet and Multi-ResNet18 with 5 corresponding image datasets as sources. In addition, we evaluate the fine-tuning performance of GPT2-small for comparison and align the parameters with the work (Zhou et al., 2023). All methods on target datasets are evaluated under the same cross-modality FSL settings (see Appendix A), with MetaSense (Gong et al., 2019) serving as the oracle baseline, trained on Conv5 with sufficient target sensor data (Table 3). For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels. As shown in Table 4 and 5, both our methods outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings. Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively. In 3-shot setting, our method achieves slightly lower accuracy than the oracle on HHAR, Gesture, and Emotion. This indicates that extremely low samples may not fully represent the test data distribution, leading to less stable feature alignment. Since the layers are reused from pre-trained sources, the performance also depends on their quality and semantic overlap with targets. Notably, the flexibility of XTransfer allows further gains in such low-shot cases by using additional source models with higher quality or closer relevance. In summary, the results show that XTransfer achieves SOTA accuracy while substantially reducing data and resource costs, providing a scalable and adaptable method for human sensing at the edge.
@@
Training resource usage. We also examine the training resource usage of XTransfer by measuring GPU utilization (Util) in the cloud. We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Chen et al., 2019), DAPN (Zhao et al., 2021), and SemiCMT (Chen et al., 2024)), excluding methods that rely solely on fine-tuning or distance-based techniques. The experimental setup remains the same as above. As shown in Table 6, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines. Notably, both Our-Single and Our-Multi achieve a substantial reduction in GPU Util by 2.6 to 3.6 times and 2.5 to 3.5 times on average, respectively, compared to baselines. This is due to our layer-wise method, which progressively repairs and recombines only a subset of layers rather than training the entire model at once. It also highlights the effectiveness of our compact connector design. Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency and scalability. Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer.
@@
MetaSense (Gong et al., 2019) 63.2 69.0 75.2 55.1 64.0 68.3 66.3 73.4 79.4 72.1 83.3 89.6 55.5 56.3 65.5 24.8 28.1 31.7
ProtoNet (Chen et al., 2019) 50.6 45.7 49.8 43.9 49.3 52.0 50.8 55.8 59.2 69.5 78.7 73.4 51.3 49.2 54.8 21.7 23.4 24.5
DAPN (Zhao et al., 2021) 48.7 51.2 54.7 56.9 61.2 63.1 45.8 49.4 50.5 75.1 78.6 79.6 48.0 50.0 58.7 24.0 24.4 25.5
MAML (Chen et al., 2019) 36.6 42.3 42.4 39.6 57.9 65.6 37.2 41.3 39.4 31.6 39.7 35.2 22.3 26.9 31.1 22.5 20.4 21.9
SemiCMT (Chen et al., 2024) 33.3 38.9 48.9 37.3 40.0 44.4 23.5 34.4 42.5 27.3 38.6 48.0 24.0 33.6 46.7 25.3 25.6 28.0
GPT2 (Zhou et al., 2023) 41.0 34.0 45.0 48.0 35.0 36.0 49.0 54.0 63.0 52.0 71.0 75.0 - - - - - -
@@
ProtoNet (Chen et al., 2019) 0.51 0.46 0.50 0.44 0.49 0.52 0.51 0.56 0.59 0.69 0.79 0.73 0.51 0.49 0.55 0.22 0.23 0.25
DAPN (Zhao et al., 2021) 0.49 0.51 0.55 0.57 0.61 0.63 0.46 0.49 0.51 0.75 0.79 0.80 0.48 0.50 0.59 0.24 0.24 0.26
MAML (Chen et al., 2019) 0.37 0.42 0.42 0.40 0.58 0.66 0.37 0.41 0.39 0.32 0.40 0.35 0.22 0.27 0.31 0.22 0.20 0.22
SemiCMT (Chen et al., 2024) 0.33 0.39 0.49 0.37 0.40 0.44 0.24 0.34 0.43 0.27 0.39 0.48 0.24 0.34 0.47 0.25 0.26 0.28
GPT2 (Zhou et al., 2023) 0.11 0.09 0.12 0.13 0.09 0.09 0.13 0.14 0.17 0.14 0.19 0.20 - - - - - -
@@
DAPN (Zhao et al., 2021) 48.85 49.19 51.79 44.62 49.49 48.81 42.73 48.71 50.59 43.88 56.63 57.51 44.67 59.73 63.12 46.66 47.32 48.48
MAML (Chen et al., 2019) 59.59 79.32 89.72 52.58 70.27 77.52 53.11 66.75 73.21 72.56 85.47 92.22 63.39 72.78 75.01 44.52 55.46 61.36
SemiCMT (Chen et al., 2024) 55.60 73.43 78.91 38.83 50.33 67.89 57.02 72.16 98.04 75.14 79.12 81.79 60.30 63.43 68.60 41.54 50.05 63.43
@@ 7 Conclusion
This paper presents XTransfer, a novel method that enables modality-agnostic few-shot model transfer with resource-efficient design for human sensing applications. XTransfer harnesses the power of pre-trained models as free sources and requires only few sensor data to create compact models. It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a scalable and adaptable method for advancing human sensing.
@@ A.1 Settings for learning with few data
A typical FSL method consists of two stages–meta-training and meta-testing (Chen et al., 2019). In meta-training stage, FSL requires a large-scale source dataset from the same modality (i.e., typically specific to the application) to train source models from scratch. The source dataset is divided into a support set and a query set  [^4] The support set and query set are assumed to be drawn from the same data distribution (Triantafillou et al., 2020).. Data for each set is loaded according to the n-way (i.e., number of classes) and n-shot (i.e., number of samples per class) format (Song et al., 2023). The meta-testing stage follows the same structure but uses unseen target data. Specifically, the target support set is used to adapt the source models, which are then evaluated on the target query set. Leave-One-Out Cross-Validation (LOOCV) is widely used as the de facto standard for evaluating datasets involving multiple users in human sensing (Gong et al., 2019). LOOCV works by leaving one user out as the target for meta-testing while using the remaining users to train the source models during meta-training in each round. The MetaSense (Gong et al., 2019) is used as an oracle baseline under FSL settings, where sufficient target sensing datasets are provided during meta-training stage. Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same modality (Oh et al., 2022). In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but extend it to transfer pre-trained source models across different modalities.
@@ A.2 Setups in preliminary study
As shown in Figure 2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ (López et al., 2022). In Figure 2(b), we test 3 source models (using a pre-trained ResNet18 from PyTorch Hub (Hub, 2022)) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Table 2. We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Table 2. We use a default 5-shot cross-modality FSL setting with LOOCV enabled.
@@ A.3 Setups in layer-wise analysis
In Section 3, we conduct a layer-wise analysis to gain insights into cross-modality model transfer with few sensor data. To set up, we continue to use the standard ResNet18 (Hub, 2022) and segment it as 9 L-Blocks, as shown in Table 3. We also employ the default reshaping (see Section B.1) on HHAR as target with 5-shot setting, and use the original source model input (miniImageNet) as reference, marked as Original. We examine the single-source baselines, including Transfer Learning (TL) (Jiang et al., 2022), Fine-tuning (FT) (Oh et al., 2022), MAML (Chen et al., 2019), DAPN (Zhao et al., 2021), and MetaSense (Gong et al., 2019) as the oracle baseline, shown in Table 1. To precisely track the MMC shift changes at each layer during fine-tuning,
@@ A.4 Layer-wise metric
To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) (Shahapure & Nicholas, 2020) as the layer-wise metric to evaluate how well the features are being fine-tuned. The S-score is in the range from -1 to 1 and being better if closing to 1. To verify whether the S-score indicates the layer performance, we observe the changes in the S-score and accuracy at each layer using the TL (Jiang et al., 2022) and Original methods. Figure 6(a) also shows that S-score and accuracy by Original are positively correlated, converging in a layer-wise manner. TL’s S-score also performs equivalently to its accuracy (i.e., S-score drop along with accuracy drop) in detecting the negative transfer across different layers.
@@ A.5 Implementation
We implement XTransfer using PyTorch version 1.4.0. We also benchmark XTransfer with a full development cycle of cross-modality FSL tasks, including pre- and post-processing, training, and inference on commercial edge devices. Once pre-trained source models and few sensor data are processed in the cloud via XTransfer, compact models can be seamlessly converted to various DL platforms (e.g., PyTorch Mobile, TFLite) for broader edge deployments. To enable reproducibility of the results, we fix all random seeds in the code, set the evaluation mode consistently when evaluating output models, and enable CUDA convolution determinism. To measure model resource overhead (e.g., FLOPs and parameters), we use the standard PyTorch OpCounter. For regression tasks, class-based transformation (Shi et al., 2022) enables compatibility.